the perimeter of a rectangular garden is 56ft. The length is 4 more ft than the width. What are the dimensions of the garden?
the perimeter of a rectangular garden is 56ft. The length is 4 more ft than the width. What are the dimensions of the garden?
P=56
Equation:p=2L+2W
L=w+4
56 = 2(w+4)+2w
56 = 2w+8+2w
56 = 4w+8
56-8=4w+8-8
48=4w
48/4=4w/4
12=w is the width
Plug in equation above to get the length:
L=12+4=16

Answer by chiexpert(48)

the perimeter of a rectangular garden is 56ft. The length is 4 more ft than the width. What are the dimensions of the garden?
P=56
Equation:p=2L+2W
L=w+4
56 = 2(w+4)+2w
56 = 2w+8+2w
56 = 4w+8
56-8=4w+8-8
48=4w
48/4=4w/4
12=w is the width
Plug in equation above to get the length:
L=12+4=16